. The measure of the length of a field is 2x + 3 ft and the measure of the width of a field is x + 10 ft. Write an expression for its perimeter
Answer by mananth(16946)

length = 2x+3
width = x+10
...
Perimeter = 2(2x+3+x+10)
P= 2(3x+13)
P=6x+26
...
m.ananth@hotmail.ca